# iTFKAN: Interpretable Time Series Forecasting with Kolmogorov–Arnold Network

### v1

@@ 5 Interpretability Analysis
$\ell_{\mathrm{reg}}=\sum_{d=1}^{D}\sum_{j=1}^{J}\sum_{i=1}^{I}||\phi^{d}_{i,j}||_{2},~{}||\phi||_{2}=\frac{1}{O}\sum_{o=1}^{O}a_{o}^{2},$ (13)
where $||\phi||_{2}$ denote the L2 norm of a univariate activation function $\phi$. Thus, the loss function of iTFKAN is calculated as follows:
@@
After sparsification, we prune the full trained KAN networks into smaller sub-networks to highlight the most important parts. Considering that each time data point of input $\mathbf{X}_{t}$ plays an important role in time-series prediction, we prune on the edges rather than the network nodes. After training, the activation function with a high L2 norm plays a significant role in the KAN network. Therefore, in line with the sparsification criterion, a univariate activation function $\phi$ will be pruned if its L2 norm $||\phi||_{2}$ is less than a threshold $\tau$.
@@ 6.3 Interpretability Case Studies
Technically, we set the threshold $\tau=5e-4$ and retain the top three significant edges for a visualization based on the L2 norm $||\phi||_{2}$ of the univariate activation function. These edges emphasize the data flow between different layers. We highlight an important node in the first layer and its connecting nodes in subsequent layers for further analysis.
@@
Figure 2: The symbolic results and data flow for the SeasonalKAN module of the iTFKAN model trained on ETTh1 dataset. ‘HUFL’ denotes the variable selected for analysis in ETTh1. The red markers label the important inputs and outputs associated with the specified node in layer one. The width of the links is proportional to the L2 norm $||\phi||_{2}$ of the edge, displaying the importance of the edges.

### v2

@@ 5 Interpretability Analysis
[…]
where $\phi||_{2}$ denote the L2 norm of a univariate activation function $\phi$. Thus, the loss function of iTFKAN is calculated as follows:
@@
After sparsification, we prune the full trained KAN networks into smaller sub-networks to highlight the most important parts. Considering that each time data point of input $\mathbf{X}_{t}$ plays an important role in time-series prediction, we prune on the edges rather than the network nodes. After training, the activation function with a high L2 norm plays a significant role in the KAN network. Therefore, in line with the sparsification criterion, a univariate activation function $\phi$ will be pruned if its L2 norm $\phi||_{2}$ is less than a threshold $\tau$.
@@ 6.3 Interpretability Case Studies
Technically, we set the threshold $\tau=5e-4$ and retain the top three significant edges for a visualization based on the L2 norm $\phi||_{2}$ of the univariate activation function. These edges emphasize the data flow between different layers. We highlight an important node in the first layer and its connecting nodes in subsequent layers for further analysis.
@@
Figure 2: The symbolic results and data flow for the SeasonalKAN module of the iTFKAN model trained on ETTh1 dataset. ‘HUFL’ denotes the variable selected for analysis in ETTh1. The red markers label the important inputs and outputs associated with the specified node in layer one. The width of the links is proportional to the L2 norm $\phi||_{2}$ of the edge, displaying the importance of the edges.